40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2026     Entry time: Wed Sep 30 01:04:56 2009
Author: rob 
Type: Update 
Category: Computers 
Subject: grief 

much grief.  somehow a burt restore of c1iscepics failed to work, and so the LSC XYCOM settings were not correct.  This meant that the LSC whitening filter states were not being correctly set and reported, making it difficult to lock for at least the last week or so.

Attachment 1: C1LSC_XYCOM.jpg  128 kB  | Hide | Hide all
C1LSC_XYCOM.jpg
ELOG V3.1.3-